OpenStudy (anonymous):

brandy and adriana work at a child care center. together they cared for 32 children this week. brandy cared for 0.6 times as many children as adriana.how many children did they care for each?

OpenStudy (anonymous):

help..

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

! !

OpenStudy (anonymous):

a + b = 32 b = (0.6)(a) a + (0.6)(a) = 32 (1.6)(a) = 32 a = 20 So, b = 12

OpenStudy (anonymous):

Good luck to you in all of your studies and thx for the recognition! @Fireflame13

OpenStudy (anonymous):

you saved my grade. thank you. fan!

OpenStudy (anonymous):

You're quite welcome and nice working with you!